OpenStudy (anonymous):

Five years ago my grandfather was five times as old as i was. Three years from now my grandfather will be three times as old as i will be. How old am i now?

OpenStudy (anonymous):

Use this instead: 5(x) = 3 + 3x 5x = 3 + 3x = 8x = 3 8x/8 = 3/8 x = .375 You're 0.375 years old.

OpenStudy (anonymous):

how tht makes no sense

hero (hero):

lol wth

hero (hero):

total complete fail

OpenStudy (anonymous):

You are really young >_<

OpenStudy (anonymous):

me

hero (hero):

You're younger

hero (hero):

matthew, you can't be serious

OpenStudy (anonymous):

matthewbacchus, rough night eh?

OpenStudy (anonymous):

right answer anyone?

OpenStudy (anonymous):

lol stay on topic guize

OpenStudy (anonymous):

lol u guys r so cute

OpenStudy (anonymous):

i'll try to , but if you keep making such horrendous mathematical errors it hard to do so

OpenStudy (anonymous):

Instead why don't you tell me and Maniqueen123 the proper steps pl0x

OpenStudy (anonymous):

lol yea i stil have like 30 more math problems after this one lol like literally

OpenStudy (anonymous):

I think it has to do with Functions, I failed badly with those...hehe

OpenStudy (anonymous):

5x-5=3+3x 2x=8 x=4 years old

OpenStudy (anonymous):

:) its ok i suk at math in general

OpenStudy (anonymous):

wait. i fail :D

OpenStudy (anonymous):

thnx soo much!!!

OpenStudy (anonymous):

+1 for you tcar2975

OpenStudy (anonymous):

rlly ? wait so is it wrong or ri?

OpenStudy (anonymous):

5 years ago. would be -1. i failed. haha. wrong setup i guess :/

OpenStudy (anonymous):

-1 for you tcar2975

OpenStudy (anonymous):

just wake me up wen someone gets the rite answer

OpenStudy (anonymous):

you must verify by replying with proper equation or steps pl0z

OpenStudy (anonymous):

hero u there

OpenStudy (anonymous):

5x = 3x + 3

OpenStudy (anonymous):

i think

OpenStudy (anonymous):

oh oh 5x = 3 + 3x 2x = 3 x = 1 1/2 years old

OpenStudy (anonymous):

wat ev who cares thnx neways :)

OpenStudy (anonymous):

NO you have a sit over there and wait this out, you might never know how to solve this properly!

OpenStudy (rhystic):

hmm are you sure thats worded correctly? if you set up the problem right its not possible I think